OpenStudy (anonymous):

Find the sum. 4x^2 + 7x^2 =

OpenStudy (anonymous):

\[\Large 4x^2+7x^2=x^2(4+7)=11x^2\]

OpenStudy (anonymous):

YEAH 11X^2